MIDC/Data Quality Statements
Total Records: 30, displaying 1-25
[Next Records]
R#SiteSerial NumberMeasurementStart DateStart TimeEnd DateEnd TimeData Quality Statement
1BMSPhotometer TWC12/28/201306:46:0012/28/201315:59:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
2BMSPhotometer TWC12/21/201307:00:0012/21/201316:22:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
3BMSPhotometer TWC12/21/201307:00:0012/21/201316:22:00Due to a scheduled NREL power outage, to upgrade campus electrical infrastructure, The following BMS instruments were not pointing at the sun during this time: Direct NIP #1, LI-201, LI-190, NIP RG780, TWC; and the following instruments were un-shaded: Diffuse CM22, CUV4, CG4. The Middle Sci-Tec tracker for these instruments were temporarily put on a secondary UPS which failed immediately.
4BMSPhotometer TWC09/13/201316:08:0009/14/201309:55:00A GFCI on the North tracker post was tripped due to water logged ventilators (caused by rain showers). Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
5BMSPhotometer TWC07/13/201316:23:0007/14/201308:29:00A GFCI on the North tracker post was tripped due to water logged ventilators (caused by rain showers). Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
6BMSPhotometer TWC05/01/201313:27:0005/02/201306:40:00A GFCI on the North tracker post was tripped due to thunderstorms. Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
7BMSPhotometer TWC03/02/201313:06:0003/02/201318:24:00Due to a scheduled NREL power outage, to upgrade campus power capacity, The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, NIP #1, LI-201, NIP RG780, CUVA2, CUVB2, TWC; and the following instruments were un-shaded: Diffuse 8-48, PSP, CM22, CUV4, PIR, CG4. The Sci-Tec tracker for these insruments are protected by UPS, however the battery depleted at 13:06 MST (the power outage started at 8:06 MST)
8BMSPhotometer TWC03/02/201308:06:0003/02/201318:24:00Due to a scheduled NREL power outage, to upgrade campus power capacity, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
9BMSPhotometer TWC07/06/201219:35:0007/09/201210:18:00A GFCI on the North tracker post was tripped due to thunderstorms. Data from the following instruments are bogus due to not having power during this time: CUVA2, CUVB2, TWC, PGS-100, POM-01.
10BMSPhotometer TWC05/26/201206:15:0005/26/201207:12:00Due to a planned power outage for SRRL electrical work, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, PGS-100, POM-01, TWC.
11BMSPhotometer TWC11/12/201111:59:0011/12/201113:26:00Due to an undervoltage followed by a power outage at 13:12, the following BMS instruments may be reading bogus or invalid data: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TWC.
12BMSPhotometer TWC09/08/201105:56:0009/08/201109:38:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, PGS-100, POM-01, TWC.
13BMSPhotometer TWC07/12/201122:57:0007/13/201107:40:00Due to a tripped GFCI outlet at SRRL the following instruments were without power during this time: TWC & POM-01 Photometers, Direct CUVA2 & CUVB2, and the PGS100. The readings are either zero or bogus and should not be used. GFCI was tripped due to Eppley venitlators which were water logged by thunderstorm.
14BMSPhotometer TWC06/14/201113:01:0006/14/201113:14:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
15BMSPhotometer TWC05/25/201105:19:0005/26/201112:16:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
16BMSPhotometer TWC05/17/201114:20:0005/17/201114:48:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
17BMSPhotometer TWC05/07/201105:55:0005/09/201107:58:00There was a planned power outage for SRRL construction activities that started at 5:18 MST. Unfortunately, the backup generator provided to SRRL was providing bad power to the UPS powering the Sci-Tec trackers, therefore the UPS failed at 5:55 MST. Unfortunately, when normal AC power was restored at 11:45 MST, the Sci-Tec trackers had lost all their configuration due to UPS cycling on and off during the outage, therefore were not tracking. The tracker settings were restored on May 9 at 7:58 MST. The following BMS instruments were not pointing at the sun during this time: Direct CH1, NIP #2, NIP #1, LI-201, NIP RG780, CUVA2, CUVB2, TWC; and the following instruments were un-shaded: Diffuse 8-48, PSP, CM22, CUV4, PIR, CG4
18BMSPhotometer TWC05/07/201105:18:0005/07/201111:45:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
19BMSPhotometer TWC04/15/201106:11:0004/15/201106:50:00Due to a planned power outage for SRRL construction activities, the following BMS instruments were down: TSP-1, TSP-700, UVSAT, UVSBT, CUVA1, CUVA2, CUVB1, CUVB2, UVB-1, 501A, MS210W, SPN1, TSI-880, PGS-100, POM-01, TWC, EFM-100.
20BMSPhotometer TWC03/18/201100:00:0003/18/201107:57:00The Sci-Tec #2 tracker was jammed on the shading plate. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201, TWC Photometer.
21BMS0001Photometer TWC10/30/201000:00:0011/01/201011:27:00The Sci-Tec #2 tracker was apparently not working during this time. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201, TWC.
22BMSPhotometer TWC12/09/200912:11:0012/09/200914:22:00Due to a power outage at NREL, followed by bad power, the SRRL tracker UPS shutdown. The following BMS instruments were not pointing at the sun: Direct CH1, NIP1, NIP2, NIP-RG780, LI201, TWC500, CUV4; and the following instruments were unshaded: 8-48, CM22, PSP, CUV4, PIR, CG4.
23BMS0001Photometer TWC07/30/200912:45:0007/30/200923:59:00The Sci-Tec #2 tracker was apparently not working during this time. The following instruments were unshaded: CM22, CUV4, CG4; and the following instruments were not pointing at the sun: NIP #1, NIP RG780, LI-201.
24BMSPhotometer TWC11/15/200814:02:0011/15/200814:32:00Due to a planned power outage at SRRL, the following BMS instruments were not pointing at the sun: Direct CH1, NIP1, NIP2, LI201, TWC500, CUV4.
25BMS0001Photometer TWC09/30/200814:18:0010/01/200815:00:00The TWC photometer was not properly aligned with the sun after being relocated to the Sci-Tec tracker. Shims were added on 10/1/08 to improve alignment.